A ball of mass 5 kg is dropped from a height of 1.4 m (from the ground). ) onto a massless spring (a spring with a length balance of 0.5 m). The ball compresses the spring a distance of 0.29 m while at rest. Calculation springs.


1
Expert's answer
2021-11-11T17:08:54-0500

h = height fallen by ball

"\frac{1}{2}kx^2 = mgh \\

k = \frac{2mgh}{x^2} \\

h = (1.4-0.5)+0.29 = 1.19 \; m \\

x = 0.29 \; m \\

k = \frac{2 \times 5 \times 9.81 \times 1.19}{(0.29)^2} = 1388 \;N"

Answer: 1388 N